# An extrapolated and provably convergent algorithm for nonlinear matrix decomposition with the ReLU function

### v1

@@ 1. Step 1
    We first prove that there exists a range of values larger than 1 for $\alpha$ such that the scheme (21) decreases the residual, that is,
    (25) $\lVert S\rVert_{F}^{2}-\lVert S(\alpha)\rVert_{F}^{2}>0.$
    Specifically, we split the residual reduction in (25) into the contribution given by the update of each variable separately. In particular, Lemma 4.5 characterizes the residual reduction after the update of $W$ and $H$, showing that it is nonnegative. Lemma 4.6 and Lemma 4.7 describe the residual reduction after updating the entries of $Z$.
@@ 3. Step 3
    We use the residual reduction that we characterized in Step 1 to prove subsequence convergence of the eBCD-NMD algorithm in Theorem 4.9, showing that a subsequence of the eBCD-NMD iterates satisfies the KKT conditions at the limit. This is done analyzing all possible situations that the restarting scheme arises.
@@
Let $U$ be an orthogonal basis of $\mathcal{R}(H^{T})$. If $H$ is full rank, we can define the orthogonal projections onto $\mathcal{R}(W(\alpha))$ and $\mathcal{R}(H^{T})$ as:
(26) $E(\alpha)=W(\alpha)W(\alpha)^{T}=Q(\alpha)Q(\alpha)^{T},$
$P=UU^{T}=H^{T}(HH^{T})^{-1}H.$
@@
Therefore, after the first two steps, the residual reduction is strictly positive unless $W(\alpha)H(\alpha)=WH$. We look at the residual reduction after updating the entries of $Z$ in $\Omega^{C}$ (Lemma 4.6) and in $\Omega$ (Lemma 4.7). In the following lemmas, we assume that there exists an interval $[1,\widehat{\alpha}]$, such that $Z_{\alpha}H^{T}$ and $ZH^{T}$ have the same rank. This hypothesis is needed to ensure the continuity of the residual reduction $\sigma_{Z}(\alpha)$ in a range of $\alpha>1$; more details are given in the Appendix.
@@
Let $(W(\alpha),H(\alpha))$ be generated by the scheme in (21). Furthermore, assume there exists $\widehat{\alpha}>1$ such that $\operatorname{rank}(Z_{\alpha}H^{T})=\operatorname{rank}(ZH^{T})$ for $\alpha\in[1,\widehat{\alpha}]$, then there exists $\alpha_{*}>1$ such that for every $\alpha\in(1,\min(\alpha_{*},\widehat{\alpha})]$,
@@
A similar result is proved in [43] for matrix completion. In their less restrictive setting, the residual reduction after updating the entries of $Z$ in $\Omega^{C}$ is independent of $\alpha$ and nonnegative. However, due to the additional constraint $P_{\Omega^{C}}(Z)\leq 0$ in the 3B-ReLU-NMD formulation, the residual reduction in (28) of the eBCD-NMD algorithm is only locally nonnegative in a range where $\alpha$ is larger than one.
@@
(30) $\lVert S\rVert_{F}^{2}-\lVert S(\alpha)\rVert_{F}^{2}=\sigma_{WH}(\alpha)+\sigma_{Z}(\alpha)>0\quad\forall\alpha\in[1,\min(\widehat{\alpha},\tilde{\alpha})].$
Corollary 4.8 guarantees that there exist some values of the extrapolation parameter $\alpha$ strictly larger than 1 resulting in a positive residual reduction for eBCD-NMD (21).
@@
Primal feasibility: $P_{\Omega}(Z-X)=0,$
$P_{\Omega^{C}}(Z)\leq 0,$
@@
where the multiplier matrices $\Lambda$ and $\Omega$ measure the residual in $\Omega$ and $\Omega^{C}$ respectively, that is, $\Lambda=P_{\Omega}(WH-Z)$ and $\Sigma=P_{\Omega^{C}}(\max(0,WH))$. Our goal is to prove that, under some mild assumptions, a subsequence of the sequence generated by eBCD-NMD satisfies the KKT condition at the limit. Specifically, one can verify that the eBCD-NMD sequence satisfies the complementary slackness condition and the stationarity condition for $Z$ at any point. Therefore, we just need to show that the optimality conditions for $W$ and $H$ are satisfied by a subsequence of the eBCD-NMD scheme at the limit, that is,
@@
Let $(W^{k+1},H^{k+1},Z^{k+1})$ be generated by eBCD-NMD (Algorithm 2), and assume that $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ is bounded for all $k$. Then there exists a bounded subsequence of $(W^{k+1},H^{k+1},Z^{k+1})$ whose limit point satisfies the KKT conditions (32).
@@
We first discuss the boundness of the two sequences $\{W^{k+1}\}$ and $\{H^{k+1}\}$ produced by Algorithm 2. Te boundness of $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ implies that $\{Z^{k}\}$ and $\{W^{k}H^{k}\}$ are bounded. This also means that $\{Z_{\alpha_{k}}\}$ is a bounded sequence by construction. Moreover, the sequence $\{W^{k+1}\}$ contains orthogonal matrices whose columns are orthogonal bases for $\mathcal{R}(Z_{\alpha_{k}}(H^{k})^{T})$, thus it is bounded. Finally, $H^{k+1}=(W^{k+1})^{T}Z_{\alpha_{k}}$ and since $\{W^{k+1}\}$ and $\{Z_{\alpha_{k}}\}$ are bounded, then also $\{H^{k+1}\}$ is bounded.
@@
where $I$ is the identity matrix of dimension equal to the number of columns of $Q^{k+1}$ or equivalently to the rank of $Z_{\alpha_{k}}(H^{k})^{T}$.
Next, we define the set of the iteration indices where the update of $Z^{k}$ results in a positive residual reduction. Let =$\mathcal{F}=\left\{k\ :\ \sigma_{Z}(\alpha_{k})\geq 0\right\}$ and $\mathcal{F}^{C}$ its complement. Therefore, denoting $E^{k}=E(\alpha_{k})$, since $H^{k+1}$ is full rank, we can use Lemma 4.5 and summing over all $k\in\mathcal{F}$, we get
@@
which implies
$\lim_{k\to\infty,k\in\mathcal{F}}\lVert E^{k}S^{k}\rVert_{F}^{2}=0.$
Let us recall that $P_{\Omega}(X)=P_{\Omega}(Z^{k})$, $P_{\Omega^{C}}(Z^{k})=P_{\Omega^{C}}(\min(0,W^{k}H^{k}))$ and using (26), we have
$S^{k}P^{k}=(Z^{k}-W^{k}H^{k})P^{k}=(Z^{k}-W^{k}H^{k})U^{k}(U^{k})^{T}.\$
Since $U^{k}$ is an orthonormal basis for $\mathcal{R}((H^{k})^{T})$ and $\{H^{k}\}$ bounded,
@@
(38) $\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}\leq(\bar{\alpha})^{2}(\lVert S^{k}P^{k}\rVert_{F}^{2}+\lVert E^{k}S^{k}(I-P^{k})\rVert_{F}^{2}).$
Combining (38) and (35) we get
$\lim_{k\to\infty,k\in\mathcal{F}}\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}=0,$
thus, from (37),
@@
Case 2a: |$|\{k\in\mathcal{F}^{C}\ |\ \delta(\alpha_{k})>\bar{\delta}\}|<\infty$. For $k$ sufficiently large, we have
$\lVert S^{k+1}\rVert_{F}^{2}\leq\bar{\delta}\lVert S^{k}\rVert_{F}^{2}.$
Therefore, it holds $\lim_{k\to\infty,k\in\mathcal{F}^{C}}\lVert S^{k}\rVert_{F}^{2}=0,$ which means that the sequence converges to a global minimizer of the problem.
Case 2b: |$|\{k\in\mathcal{F}^{C}\ |\ \delta(\alpha_{k})>\bar{\delta}\}|=\infty$. The eBCD-NMD set the acceleration parameter $\alpha_{k}$ to 1 for an infinite number of iterations. Denote $\mathcal{J}_{1}$ the set of indices in which $\alpha_{k}=1$, a similar relation to (34) holds:
@@
Therefore, one can replicate the proof for Case 1 and the subsequence whose indexes are in $\mathcal{J}_{1}$ satisfies the optimality conditions at the limit.
We emphasize that, similarly to BCD-NMD, we need a strong assumption, namely the boundness of $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ which guarantees that the eBCD-NMD sequence remains bounded. This hypothesis serves the same purpose as the existence of the limit point in the BCD-NMD algorithm, since it allows us to exclude pathological cases where the infimum of the problem is not attained.
@@ 5 Numerical experiments
In this section, we present numerical experiments on synthetic and real-world data sets, for different applications of ReLU decomposition: matrix completion with ReLU sampling, the recovery of Euclidean distance matrices, low-dimensional embedding on text data, and the compression sparse dictionaries. Our main purpose is to show that eBCD-NMD behaves favorably compared to the state of the art, namely Naive-NMD [34], EM-NMD [34], A-NMD [35], and 3B-NMD [35], while having convergence guarantees.
In all experiments, we use random initialization: entries of $W$ and $H$ are sampled from a Gaussian distribution, W = randn(m,r) and H = randn(r,n) in MATLAB. We then scale them so that the initial point matches the magnitude of the original matrix, that is, we multiply $W$ by $\sqrt{\lVert X\rVert_{F}}/\lVert W\rVert_{F}$ and $H$ by $\sqrt{\lVert X\rVert_{F}}/\lVert H\rVert_{F}$. If not otherwise specified, we stop each algorithm when we reach a small relative residual, namely $\Gamma_{k}=\lVert Z^{k}-W^{k}H^{k}\rVert_{F}/\lVert X\rVert_{F}\leq 10^{-9}$, or a fixed time limit is reached. The time limit will be selected depending on the size of the data. For the algorithms solving Latent-ReLU-NMD, the low-rank product $W^{k}H^{k}$ in $\Gamma_{k}$ is substituted with the matrix $\Theta^{k}$. The first stopping criteria can be achieved only if the target matrix admits an (almost) exact ReLU decomposition. The algorithms are implemented in MATLAB R2021b on a 64-bit Samsung/Galaxy with 11th Gen Intel(R) Core(TM) i5-1135G7 @ 2.40GHz and 8 GB of RAM, under Windows 11 version 23H2. The code is available online from https://github.com/giovanniseraghiti/ReLU-NMD .
The parameters of eBCD-NMD are the ones specified in Algorithm 2. For the parameters of the other methods, we use the setting from [35].
@@ 5.1 Matrix completion with ReLU sampling
In this subsection, we consider the same experiment used in [26] for symmetric matrix completion with ReLU sampling, adapted to the non-symmetric case. Specifically, we randomly generate the entries $W$ and $H$ from a Gaussian distribution, as for our initializations, and we construct $\Theta^{t}=WH$, and set $X=\max(0,\Theta^{t})$ in the noiseless case and $\max(0,\Theta^{t}+N)$ in the noisy case, where $N=\sigma*\tilde{N}*\lVert\Theta^{t}\rVert_{F}/\lVert\tilde{N}\rVert_{F}$, $\tilde{N}=$ randn(m,n), and $\sigma\in(0,1)$. In our experiment, we set $m=n=1000$, $r=20$ and $\sigma=10^{-2}$. Note that since $W$, $H$ and $N$ are Gaussian, $X$ is approximately $50\%$ sparse. We present the loglog plot of the average residual $\Gamma_{k}$ against the average CPU time over 20 different randomly generated matrices $X$ in Figure 2. The algorithms are stopped when the objective functions in (2) and (3) achieve a relative residual $\Gamma_{k}$ below $10^{-9}$ in the noiseless case, or below $\sigma=10^{-2}$ in the noisy case. We also report the average CPU time to reach these values in Table 1.
@@
BCD-NMD eBCD-NMD 3B-NMD Naive-NMD A-NMD EM-NMD
Noiseless 7.92 $\pm$ 0.16 0.90 $\pm$ 0.04 1.95 $\pm$ 0.04 22.0 $\pm$ 0.78 6.37 $\pm$ 0.33 26.7 $\pm$ 0.70
Noisy 1.14 $\pm$ 0.05 0.19 $\pm$ 0.01 0.66 $\pm$ 0.02 3.17 $\pm$ 0.12 1.54 $\pm$ 0.19 2.70 $\pm$ 0.15
Table 1: Average CPU time in seconds and standard deviation to reach a tolerance on the relative residual $\Gamma_{k}$ of $10^{-9}$ in the noiseless case, and of $10^{-2}$ in the noisy case. Best values are highlighted in bold.
Figure 2 shows that all the algorithms reach the tolerance on the relative residual. Moreover, eBCD-NMD outperforms all the other methods, being considerably faster both in the noiseless and noisy cases, as Table 1 confirms. In fact, compared to the second best, eBCD-NMD is on average more than twice faster in the noiseless case, and more than three time faster in the noisy case (each time, the second best is 3B-NMD).

### v2

@@ 1. Step 1
    We first prove that there exists a range of values larger than 1 for $\alpha$ such that the scheme (20) decreases the residual, that is, $\lVert S\rVert_{F}^{2}-\lVert S(\alpha)\rVert_{F}^{2}>0$. Specifically, we split this residual reduction into the contribution given by the update of each variable separately: Lemma 4.5 characterizes the residual reduction after the update of $W$ and $H$, showing that it is nonnegative. Lemma 4.6 and Lemma 4.7 describe the residual reduction after updating the entries of $Z$.
@@ 3. Step 3
    We use the residual reduction that we characterized in Step 1 to prove the subsequence convergence of the eBCD algorithm in Theorem 4.9, showing that a subsequence of the eBCD iterates satisfies the KKT conditions at the limit. This is done by analyzing all possible situations that the restarting scheme arises.
@@
Let $U$ be an orthonormal basis of $\mathcal{R}(H^{T})$. If $H$ is full rank, we can define the orthogonal projections onto $\mathcal{R}(W(\alpha))$ and $\mathcal{R}(H^{T})$ as:
(24) $E(\alpha)=W(\alpha)W(\alpha)^{T}=Q(\alpha)Q(\alpha)^{T},\qquad P=UU^{T}=H^{T}(HH^{T})^{-1}H.$
@@
Therefore, after the first two steps, the residual reduction is strictly positive unless $W(\alpha)H(\alpha)=WH$. We look at the residual reduction after updating the entries of $Z$ in $\Omega^{C}$ (Lemma 4.6) and in $\Omega$ (Lemma 4.7). In the following lemmas, we assume that there exists an interval $1,\widehat{\alpha}$], such that $Z_{\alpha}H^{T}$ and $ZH^{T}$ have the same rank. This hypothesis is needed to ensure the continuity of the residual reduction $\sigma_{Z}(\alpha)$ in a range of $\alpha>1$; more details are given in the Appendix.
@@
[…]
@@
A similar result is proved in [wen2012solving] for matrix completion. In their less restrictive setting, the residual reduction after updating the entries of $Z$ in $\Omega^{C}$ is independent of $\alpha$ and nonnegative. However, due to the additional constraint $P_{\Omega^{C}}(Z)\leq 0$ in the 3B-RMD formulation, the residual reduction in (26) of the eBCD algorithm is only locally nonnegative in a range where $\alpha$ is larger than one.
@@
[…]
Corollary 4.8 guarantees that there exist some values of the extrapolation parameter $\alpha$ strictly larger than 1 resulting in a positive residual reduction for eBCD (20).
@@
Primal feasibility: $P_{\Omega}(Z-X)=0,P_{\Omega^{C}}(Z)\leq 0,$
@@
where the multiplier matrices $\Lambda$ and $\Omega$ measure the residual in $\Omega$ and $\Omega^{C}$ respectively, that is, $\Lambda=P_{\Omega}(WH-Z)$ and $\Sigma=P_{\Omega^{C}}(\max(0,WH))$. Our goal is to prove that, under some mild assumptions, a subsequence of the sequence generated by eBCD satisfies the KKT condition at the limit. Specifically, one can verify that the eBCD sequence satisfies the complementary slackness condition and the stationarity condition for $Z$ at any point. Therefore, we just need to show that the optimality conditions for $W$ and $H$ are satisfied by a subsequence of the eBCD scheme at the limit, that is,
@@
Let $(W^{k+1},H^{k+1},Z^{k+1})$ be generated by eBCD (Algorithm 2), and assume that $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ is bounded for all $k$. Then there exists a bounded subsequence of
$(W^{k+1},H^{k+1},Z^{k+1})$ whose limit point satisfies the KKT conditions (30).
@@
We first discuss the boundness of the two sequences $\{W^{k+1}\}$ and $\{H^{k+1}\}$ produced by Algorithm 2. Te boundness of $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ implies that $\{Z^{k}\}$ and $\{W^{k}H^{k}\}$ are bounded. This also means that $\{Z_{\alpha_{k}}\}$ is a bounded sequence by construction. Moreover, the sequence $\{W^{k+1}\}$ contains orthogonal matrices whose columns form an orthonormal basis for $\mathcal{R}(Z_{\alpha_{k}}(H^{k})^{T})$, thus it is bounded. Finally, $H^{k+1}=(W^{k+1})^{T}Z_{\alpha_{k}}$ and since $\{W^{k+1}\}$ and $\{Z_{\alpha_{k}}\}$ are bounded, then also $\{H^{k+1}\}$ is bounded.
@@
where $I$ is the identity matrix of dimension equal to the number of columns of $Q^{k+1}$ or equivalently, to the rank of $Z_{\alpha_{k}}(H^{k})^{T}$.
Next, we define the set of indices where the update of $Z^{k}$ results in a positive residual reduction. Let $\mathcal{F}=\left\{k\ :\ \sigma_{Z}(\alpha_{k})\geq 0\right\}$ and $\mathcal{F}^{C}$ its complement. Therefore, denoting $E^{k}=E(\alpha_{k})$, since $H^{k+1}$ is full rank, we can use Lemma 4.5, and summing over all $k\in\mathcal{F}$, we get
@@
which implies $\lim_{k\to\infty,k\in\mathcal{F}}\lVert E^{k}S^{k}\rVert_{F}^{2}=0$. Recall that $P_{\Omega^{C}}(Z^{k})=P_{\Omega^{C}}(\min(0,W^{k}H^{k}))$ and $P_{\Omega}(X)=P_{\Omega}(Z^{k})$. Using (24), we have $S^{k}P^{k}=(Z^{k}-W^{k}H^{k})P^{k}=(Z^{k}-W^{k}H^{k})U^{k}(U^{k})^{T}$. Since $U^{k}$ is an orthonormal basis for $\mathcal{R}((H^{k})^{T})$ and $\{H^{k}\}$ bounded,
@@
(36) $\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}\leq\bar{\alpha}^{2}(\lVert S^{k}P^{k}\rVert_{F}^{2}+\lVert E^{k}S^{k}(I-P^{k})\rVert_{F}^{2}).$
Combining (36) and (33) we get $\lim_{k\to\infty,k\in\mathcal{F}}\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}=0$, thus, from (35),
@@
Case 2a: $|\{k\in\mathcal{F}^{C}\ |\ \delta(\alpha_{k})>\bar{\delta}\}|<\infty$. For $k$ sufficiently large, $\lVert S^{k+1}\rVert_{F}^{2}\leq\bar{\delta}\lVert S^{k}\rVert_{F}^{2}$. Therefore, $\lim_{k\to\infty,k\in\mathcal{F}^{C}}\lVert S^{k}\rVert_{F}^{2}=0,$ which means that the sequence converges to a global minimizer of the problem.
Case 2b: $|\{k\in\mathcal{F}^{C}\ |\ \delta(\alpha_{k})>\bar{\delta}\}|=\infty$. eBCD sets the parameter $\alpha_{k}$ to 1 for an infinite number of iterations. Denote $\mathcal{J}_{1}$ the set of indices in which $\alpha_{k}=1$, a similar relation to (32) holds:
@@
Therefore, one can replicate the proof for Case 1, and the subsequence whose indices are in $\mathcal{J}_{1}$ satisfies the optimality conditions at the limit.
We emphasize that, similarly to BCD, we need a strong assumption, namely the boundness of $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ which guarantees that the eBCD sequence remains bounded. This hypothesis serves the same purpose as the existence of the limit point in the BCD algorithm, since it allows us to exclude pathological cases where the infimum of the problem is not attained.
@@ 5 Numerical experiments
In this section, we present numerical experiments on synthetic and real-world data sets, for different applications of RMD: matrix completion with ReLU sampling, the recovery of Euclidean distance matrices, low-dimensional embedding on text data, and the compression sparse dictionaries. Our main purpose is to show that eBCD behaves favorably compared to the state of the art, namely the Naive approach [saul2022nonlinear], EM [saul2022nonlinear], A-Naive [seraghiti2023accelerated], and e3B [seraghiti2023accelerated], while having convergence guarantees. In all experiments, we use random initialization: entries of $W$ and $H$ are sampled from a Gaussian distribution, W = randn(m,r) and H = randn(r,n) in MATLAB. We then scale them so that the initial point matches the magnitude of the original matrix, that is, we multiply $W$ by $\sqrt{\lVert X\rVert_{F}}/\lVert W\rVert_{F}$ and $H$ by $\sqrt{\lVert X\rVert_{F}}/\lVert H\rVert_{F}$. If not otherwise specified, we stop each algorithm when we reach a small relative residual, namely $\Gamma_{k}=\lVert Z^{k}-W^{k}H^{k}\rVert_{F}/\lVert X\rVert_{F}\leq 10^{-9}$, or a fixed time limit is reached. The time limit will be selected depending on the size of the data. For the algorithms solving Latent-RMD, the low-rank product $W^{k}H^{k}$ in $\Gamma_{k}$ is substituted with the matrix $\Theta^{k}$. The first stopping criteria can be achieved only if the target matrix admits an (almost) exact RMD. The algorithms are implemented in MATLAB R2021b on a 64-bit Samsung/Galaxy with 11th Gen Intel(R) Core(TM) i5-1135G7 @ 2.40GHz and 8 GB of RAM, under Windows 11 version 23H2. The code is available online from https://github.com/giovanniseraghiti/ReLU-NMD .
The parameters of eBCD are the ones specified as default in Algorithm 2. For the parameters of the other methods, we use the setting from [seraghiti2023accelerated].
@@ 5.1 Matrix completion with ReLU sampling
In this subsection, we consider the same experiment used in [liu2024symmetric] for symmetric matrix completion with ReLU sampling, adapted to the non-symmetric case. Specifically, we randomly generate the entries $W$ and $H$ from a Gaussian distribution, as for our initializations, and we construct $\Theta^{t}=WH$, and set $X=\max(0,\Theta^{t})$ in the noiseless case and $\max(0,\Theta^{t}+N)$ in the noisy case, where $\sigma*\tilde{N}*\lVert\Theta^{t}\rVert_{F}/\lVert\tilde{N}\rVert_{F}, N~=\tilde{N}=$ randn(m,n), and $\sigma\in(0,1)$. In our experiment, we set $m=n=1000$, $r=20$ and $\sigma=10^{-2}$. Note that since $W$, $H$ and $N$ are Gaussian, $X$ is approximately $50\%$ sparse. We present the loglog plot of the average residual $\Gamma_{k}$ against the average CPU time over 20 different randomly generated matrices $X$ in Figure 1. The algorithms are stopped when the objective functions in (2) and (3) achieve a relative residual $\Gamma_{k}$ below $10^{-9}$ in the noiseless case, or below $\sigma=10^{-2}$ in the noisy case. We also report the average CPU time and the standard deviation to reach these values in Table 1.
@@
BCD eBCD e3B Naive A-Naive EM
[…]
Noise iter 304 121 65 308 84 296
[…]
Noise iter 36 22 20 41 19 28
Table 1: Average CPU time in seconds, standard deviation, and average iteration number to reach a tolerance on the relative residual $\Gamma_{k}$ of $10^{-9}$ in the noiseless case, and of $10^{-2}$ in the noisy case. Best values are highlighted in bold.
Figure 1 shows that all the algorithms reach the tolerance on the relative residual. Moreover, eBCD outperforms all the other methods, being considerably faster both in the noiseless and noisy cases, as Table 1 confirms. In fact, compared to the second best, eBCD is on average more than twice faster in the noiseless case, and more than three time faster in the noisy case (each time, the second best is e3B). Table 1 shows also that A-Naive and e3B need less iterations to converge than eBCD. However, their computational cost per iteration is significantly larger.